The Speed of Light. One night you are out looking up at the stars and an extra-terrestrial spaceshipf flashesacross the sky. The ship is 50 meters long and is travelling at 95 percent of the speed of light. What would the ship's length be when measured from your earthbound frame of reference?

With solution


1
Expert's answer
2021-05-23T16:36:00-0400

For us, the length will contract:


"L=L_0\sqrt{1-\frac{v^2}{c^2}}=50\sqrt{1-(0.95c)^2/c^2}=\\
=50\sqrt{1-0.95^2}=15.6\text{ m}."